# ELITE: Enhanced Language-Image Toxicity Evaluation for Safety

### v2

@@ 1 Introduction
Vision Language Models (VLMs), which are composed of pre-trained Large Language Models (LLMs) and visual encoders, have been introduced to tackle complex multimodal tasks. Despite advancements in their capabilities and performance, VLMs remain vulnerable to malicious inputs, raising significant safety concerns and posing substantial challenges to their large-scale deployment in real-world applications (Zong et al., 2024; Carlini et al., 2023; Gong et al., 2023; Bommasani et al., 2022).
Concerns about the safety of VLMs, such as malicious users inducing harmful outputs, have been raised and several safety evaluation benchmarks have appeared to assess the safety of VLMs (Zong et al., 2024; Liu et al., 2025; Gu et al., 2024; Zhang et al., 2024; Wang et al., 2024b; Gong et al., 2023; Luo et al., 2024). These benchmarks are typically constructed by collecting image-text pairs that can trigger harmful responses, with safety evaluations conducted through automated methods using language models rather than humans (Zong et al., 2024; Luo et al., 2024; Liu et al., 2025; Zhang et al., 2024; Li et al., 2024). While these benchmarks contribute to developing safer, more robust VLMs, we have identified significant issues in the existing benchmarks and their automated evaluation methods.
@@
Third, existing benchmarks mainly consist of unsafe image-unsafe text pairs (i.e., unsafe-unsafe, safe-unsafe, or unsafe-safe pairs) (Gu et al., 2024; Zong et al., 2024; Liu et al., 2025). However, as demonstrated in the examples at the bottom of Fig. 1(c), harmful responses can also be induced through safe-safe pairs (Wang et al., 2024b). To address this issue, we propose various methods for inducing harmful responses from VLMs. As shown in Fig. 1(b), the ELITE benchmark incorporates four in-house generated methods, which improve coverage of all four image-text pair combinations. This enhances the diversity of the ELITE benchmark and enables a more comprehensive evaluation of VLM safety.
The ELITE evaluator is built upon the StrongREJECT evaluator (Souly et al., 2024), a rubric-based safety evaluation method that suggests the overestimation of the evaluation results for jailbreak techniques in LLMs. Unlike LLMs, where only the text prompt influences harmful responses, VLMs are influenced by the images as well. We find that input images can frequently cause the victim model to generate responses, such as simple image descriptions, that are unhelpful to users attempting to elicit harmful responses. To address this, we incorporate a toxicity score into the evaluation framework, enabling a more accurate assessment of how harmful the victim model’s responses truly are. This refinement ensures a more precise evaluation by distinguishing genuinely safe responses, such as merely descriptive responses. Additionally, we construct the ELITE benchmark, a high-quality benchmark, using the ELITE evaluator to filter out ambiguous image-text pairs. We achieve this by using the ELITE evaluator scores (defined in Section 3) to retain only explicit harmful responses while incorporating a diverse set of image-text pairs from both existing benchmarks and in-house generated image-text pairs.
@@ 2 Related Work
To address vulnerabilities and evaluate the safety of VLMs, various benchmarks have been developed. VLGuard (Zong et al., 2024) introduces a fine-grained evaluation benchmark that focuses on visual-linguistic reasoning, leveraging a taxonomy that categorizes potential safety issues. MM-SafetyBench (Liu et al., 2025) provides a comprehensive benchmark with image-text pairs across 13 safety-critical scenarios, emphasizing image-based manipulations and their impact on VLMs’ responses. MLLMGuard (Gu et al., 2024) evaluates safety across five dimensions—privacy, bias, toxicity, truthfulness, and legality—using a bilingual benchmark. Besides these benchmarks, others such as SPA-VL (Zhang et al., 2024), JailbreakV-28K (Luo et al., 2024), and SIUO (Wang et al., 2024b), contribute to evaluating vulnerabilities and enhancing the robustness of VLMs.
@@ 3.1 The ELITE Evaluator
Figure 2: Examples of safety evaluations about the victim model’s response by ELITE and StrongREJECT evaluator. $r$, $s$, $c$, and $t$ represent refused, specific, convincing, and toxicity, respectively. The ELITE evaluator can effectively evaluate utilizing the toxicity score to make more accurate judgments.
@@ 3.4 Benchmark Construction Pipeline
Table 3: ELITE evaluator score-based ASR of various VLMs across taxonomies. The upper group in the table represents proprietary models, and the lower group represents open-source models. The most vulnerable model is highlighted in bold and the second-most vulnerable with an underline.
@@ 4.1 Experiment Setup
We evaluate the effectiveness of the ELITE benchmark, consisting of 4,587 image-text pairs, across various VLMs, including GPT-4o, GPT-4o-mini, Gemini-2.0, Gemini-1.5, and open-source models. For open-source models, their original hyperparameters are used. We use GPT-4o as the ELITE evaluator to evaluate the safety of VLMs.
@@ 4.2 Metric
where $\text{ELITE evaluator score}_{i}$ represents the ELITE evaluator score of the $i$-th image-text pair and $N$ is the total number of image-text pairs.
@@ 4.3 Evaluation of the ELITE Benchmark
Table 4: Comparison of the average E-ASR when using different benchmarks. It highlights that the most effective benchmark for inducing harmful responses in bold and the second-most effective benchmark with an underline.
Model Benchmark Total Average
LLaVa-v1.5-7B VLGuard 2028 27.75
MM-SafetyBench 1680 45.06
MLLMGuard 532 27.26
ELITE benchmark (generated) 1054 69.17
ELITE benchmark 4587 63.59
LLaVa-v1.5-13B VLGuard 2028 28.40
MM-SafetyBench 1680 46.61
MLLMGuard 532 27.26
ELITE benchmark (generated) 1054 78.46
ELITE benchmark 4587 69.68
DeepSeek-VL-7B VLGuard 2028 16.40
MM-SafetyBench 1680 31.79
MLLMGuard 532 16.29
ELITE benchmark (generated) 1054 37.95
ELITE benchmark 4587 42.36
ShareGPT4V-7B VLGuard 2028 29.24
MM-SafetyBench 1680 48.81
MLLMGuard 532 23.51
ELITE benchmark (generated) 1054 68.50
ELITE benchmark 4587 67.16
@@ 4.4 Comparisons with Other Benchmarks
In this section, we demonstrate the superiority of both the ELITE benchmark and the ELITE benchmark (generated). Table 4 compares the E-ASR of LLaVa-v1.5 (7B, 13B), DeepSeek-VL (7B), and ShareGPT4V (7B) across existing benchmarks, including VLGuard (Zong et al., 2024), MM-SafetyBench (Liu et al., 2025), MLLMGuard (Gu et al., 2024), the ELITE benchmark, and the ELITE benchmark (generated). Note that we use publicly available benchmarks in this experiment.
@@ 5.1 Human Annotators and Datasets
The dataset used for human evaluation consists of a subset of image-text pairs, with approximately 90 pairs per taxonomy, totaling 963 pairs. For a fair comparison, instead of using the filtered responses, we also include responses from models that did not meet the filtering criteria among the three models (Phi-3.5-Vision, Llama-3.2-11B-Vision, and Pixtral-12B). Additionally, these pairs were primarily collected where the evaluation results differed between the ELITE evaluator and existing evaluation methods (StrongREJECT evaluator and safeguard models) and were randomly sampled across each taxonomy. To ensure diversity, we excluded image-text pairs that differed only in model responses from the human evaluation dataset.
@@ 6 Limitations
The ELITE evaluator aims to provide an accurate safety evaluation. However, as a rubric-based approach, its evaluation performance may vary depending on the capabilities of the VLM, which we acknowledge as a potential limitation.
Moreover, although the ELITE benchmark is constructed to cover a broad taxonomy and include diverse image-text pairs, certain cases that elicit harmful responses may still not be contained. For example, multi-turn jailbreak methods remain unincorporated.
@@ 7 Conclusion
In this paper, we propose the ELITE evaluator, an enhanced rubric-based evaluation method for assessing the safety of responses generated by VLMs. Furthermore, we introduce the ELITE benchmark, a high-quality safety benchmark of 4,587 image-text pairs filtered using the ELITE evaluator, including 1,054 in-house generated pairs designed to elicit harmful responses for robust safety evaluations of VLMs. Through experiments, we demonstrate the quality of the ELITE benchmark by comparing it with existing benchmarks and show that the ELITE evaluator aligns more closely with human judgments than existing safety evaluation methods. Our experiments highlight the need for overall improvement in safety alignment, and through this work, we aim to strengthen the safety of VLMs.
@@ A.1 Benchmark using the ELITE Evaluator Score
Table 8: ELITE evaluator score of various VLMs across taxonomies. The upper group in the table represents proprietary models, and the lower group represents open-source models. Highlight the most vulnerable model in bold and the second-most vulnerable with an underline.
@@ B.2 Previous Benchmarks
SIUO (Wang et al., 2024b) 167 Human Human, GPT-4V
@@ B.5 Model Cards
Qwen2-VL-7B (Wang et al., 2024a) 7B Custom ViT with 600M params Qwen2-7B

### v3

@@ 1 Introduction
Vision Language Models (VLMs), which are composed of pre-trained Large Language Models (LLMs) and visual encoders, have been introduced to tackle complex multimodal tasks. Despite advancements in their capabilities and performance, VLMs remain vulnerable to malicious inputs, raising significant safety concerns and posing substantial challenges to their large-scale deployment in real-world applications (Zong et al., 2024; Carlini et al., 2023; Gong et al., 2023; Bommasani et al., 2021).
Concerns about the safety of VLMs, such as malicious users inducing harmful outputs, have been raised and several safety evaluation benchmarks have appeared to assess the safety of VLMs (Zong et al., 2024; Liu et al., 2025; Gu et al., 2024; Zhang et al., 2024; Wang et al., 2025; Gong et al., 2023; Luo et al., 2024). These benchmarks are typically constructed by collecting image-text pairs that can trigger harmful responses, with safety evaluations conducted through automated methods using language models rather than humans (Zong et al., 2024; Luo et al., 2024; Liu et al., 2025; Zhang et al., 2024; Li et al., 2024). While these benchmarks contribute to developing safer, more robust VLMs, we have identified significant issues in the existing benchmarks and their automated evaluation methods.
@@
Third, existing benchmarks mainly consist of unsafe image-unsafe text pairs (i.e., unsafe-unsafe, safe-unsafe, or unsafe-safe pairs) (Gu et al., 2024; Zong et al., 2024; Liu et al., 2025). However, as demonstrated in the examples at the bottom of Fig. 1(c), harmful responses can also be induced through safe-safe pairs (Wang et al., 2025). To address this issue, we propose various methods for inducing harmful responses from VLMs. As shown in Fig. 1(b), the ELITE benchmark incorporates four in-house generated methods, which improve coverage of all four image-text pair combinations. This enhances the diversity of the ELITE benchmark and enables a more comprehensive evaluation of VLM safety.
The ELITE evaluator builds on StrongREJECT (Souly et al., 2024), extending its rubric-based evaluation to vision-language tasks by incorporating toxicity scores. This addition helps distinguish genuinely harmful outputs from merely descriptive ones, which are common in VLMs due to the influence of image input. Leveraging this evaluator, we curate the ELITE benchmark by filtering out ambiguous pairs and selecting diverse, explicitly harmful image-text examples from both existing and newly generated data, enabling more reliable safety evaluation.
@@ 2 Related Work
To address vulnerabilities and evaluate the safety of VLMs, various benchmarks have been developed, following the previous safety evaluation benchmarks for LLM (Wang et al., 2023). VLGuard (Zong et al., 2024) introduces a fine-grained evaluation benchmark that focuses on visual-linguistic reasoning, leveraging a taxonomy that categorizes potential safety issues. MM-SafetyBench (Liu et al., 2025) provides a comprehensive benchmark with image-text pairs across 13 safety-critical scenarios, emphasizing image-based manipulations and their impact on VLMs’ responses. MLLMGuard (Gu et al., 2024) evaluates safety across five dimensions—privacy, bias, toxicity, truthfulness, and legality—using a bilingual benchmark. Besides these benchmarks, others such as SPA-VL (Zhang et al., 2024), JailbreakV-28K (Luo et al., 2024), and SIUO (Wang et al., 2025), contribute to evaluating vulnerabilities and enhancing the robustness of VLMs.
@@ 3.1 The ELITE Evaluator
Figure 2: Examples of safety evaluations about the victim model’s response by ELITE and StrongREJECT evaluator. $r$, $s$, $c$, and $t$ represent refused, specific, convincing, and toxicity, respectively. The ELITE evaluator can effectively evaluate by utilizing the toxicity score to make more accurate judgments.
@@ 3.4 Benchmark Construction Pipeline
Table 3: ELITE evaluator score-based ASR of various VLMs across taxonomies. The upper group in the table represents proprietary models, and the lower group represents open-source models. The most vulnerable model is highlighted in bold and the second-most vulnerable with an underline. * denotes the model used for filtering.
@@ 4.1 Experiment Setup
We evaluate the effectiveness of the ELITE benchmark, consisting of 4,587 image-text pairs, across various VLMs, including GPT-4o (OpenAI, 2024a), GPT-4o-mini (OpenAI, 2024a), Gemini-2.0 (Google DeepMind, 2024), Gemini-1.5 (Gemini Team et al., 2024), and open-source models. For open-source models, their original hyperparameters are used. We use GPT-4o as the ELITE evaluator to evaluate the safety of VLMs.
@@ 4.2 Metric
where $\text{ELITE score}_{i}$ represents the ELITE evaluator score of the $i$-th image-text pair and $N$ is the total number of image-text pairs.
@@ 4.3 Evaluation of the ELITE Benchmark
Table 4: Comparison of the average E-ASR and ASR when using different benchmarks. It highlights that the most effective benchmark for inducing harmful responses in bold and the second-most effective benchmark with an underline.
Model Benchmark Total E-ASR ASR
LLaVa-v1.5-7B VLGuard 2028 27.75 34.82
MM-SafetyBench 1680 45.06 39.67
MLLMGuard 532 27.26 36.46
ELITE benchmark (generated) 1054 69.17 70.83
ELITE benchmark 4587 63.59 68.98
LLaVa-v1.5-13B VLGuard 2028 28.40 34.00
MM-SafetyBench 1680 46.61 41.25
MLLMGuard 532 27.26 32.65
ELITE benchmark (generated) 1054 78.46 69.24
ELITE benchmark 4587 69.68 69.99
DeepSeek-VL-7B VLGuard 2028 16.40 28.59
MM-SafetyBench 1680 31.79 38.63
MLLMGuard 532 16.29 23.35
ELITE benchmark (generated) 1054 37.95 57.83
ELITE benchmark 4587 42.36 60.83
ShareGPT4V-7B VLGuard 2028 29.24 31.98
MM-SafetyBench 1680 48.81 40.89
MLLMGuard 532 23.51 30.11
ELITE benchmark (generated) 1054 68.50 66.60
ELITE benchmark 4587 67.16 69.54
@@ 4.4 Comparisons with Other Benchmarks
In this section, we demonstrate the superiority of both the ELITE benchmark and the ELITE benchmark (generated). Table 4 compares the E-ASR of LLaVa-v1.5 (7B, 13B), DeepSeek-VL (7B), ShareGPT4V (7B), Gemma-3 (4B), and InternVL2.5 (26B) across existing benchmarks, including VLGuard (Zong et al., 2024), MM-SafetyBench (Liu et al., 2025), MLLMGuard (Gu et al., 2024), the ELITE benchmark, and the ELITE benchmark (generated). Note that we use publicly available benchmarks in this experiment.
@@
Moreover, to demonstrate that the ELITE benchmark is not overly tailored to the ELITE evaluator, we present results based on the previously adopted metric, Attack Success Rate (ASR), instead of E-ASR. These results suggest that the ELITE benchmark remains general and is not excessively influenced by the use of the ELITE evaluator.
@@ 5.1 Human Annotators and Datasets
The dataset used for human evaluation consists of a subset of image-text pairs, with approximately 90 pairs per taxonomy, totaling 963 pairs. For a fair comparison, instead of using the filtered responses, we also include responses from models that did not meet the filtering criteria among the three models (Phi-3.5-Vision, Llama-3.2-11B-Vision, and Pixtral-12B). Additionally, these pairs were primarily collected where the evaluation results differed between the ELITE evaluator and existing evaluation methods (e.g., StrongREJECT evaluator and safeguard models) and were randomly sampled across each taxonomy. To ensure diversity, we excluded image-text pairs that differed only in model responses from the human evaluation dataset.
@@ 5.3 Comparison with Safeguard Models
6 Analysis on the ELITE Evaluator & Benchmark
In this section, we note that the human evaluation dataset (Sec. 5) may be biased, as it primarily consists of samples where the evaluation results differed between the ELITE evaluator and existing methods (e.g., the StrongREJECT evaluator and safeguard models).
To enable a more rigorous analysis of the ELITE evaluator & benchmark, we conduct an additional human evaluation. This additional human evaluation dataset consists of 228 samples, 110 of which were randomly selected from the ELITE benchmark (referred to as From ELITE) and 118 that were not included (i.e., filtered out, referred to as Not From ELITE). We included at least 20 samples from each taxonomy and gathered the opinions of 3 labelers per sample, with the final labeling determined by majority decision. In total, 8 labelers were recruited for this evaluation. We provided the input image, text, and the model’s response to perform the safety judgment.
Figure 5: Pearson correlation coefficients among three labelers and the ELITE evaluator based on toxicity scores over 228 samples.
6.1 Alignment of Toxicity Scores Between ELITE and Humans
We conduct a quantitative analysis to assess how well the ELITE evaluator’s toxicity scores align with human judgment. Specifically, we compute the Pearson correlation coefficients between the ELITE evaluator (GPT-4o) and three individual human labelers across 228 samples. As shown in Fig. 5, the correlation between ELITE and Human1 is 0.7274, with 0.6447 for Human2 and 0.6496 for Human3. These values indicate a strong alignment between the ELITE evaluator and each human rater.
In contrast, the correlations among human labelers are comparatively lower: 0.5992 between Human1 and Human2, 0.6079 between Human1 and Human3, and 0.5079 between Human2 and Human3. As a result, the correlation between the ELITE evaluator and individual human labelers is higher than the inter-labelers correlations, suggesting that the ELITE evaluator provides more consistent toxicity judgments despite variability among human raters.
6.2 Validation of the ELITE Benchmark with Human Evaluation
To validate the quality of the ELITE benchmark, we use an additional human evaluation dataset to compare the proportion of harmful responses between included and excluded samples. As shown in Tab. 8, 67.27% of the From ELITE samples were labeled as unsafe by human labelers, while only 11.86% of the Not From ELITE samples were judged to be unsafe. This substantial difference confirms that the ELITE evaluator effectively selects harmful cases and filters out ambiguous ones. By incorporating a toxicity score, it further enhances the precision of safety evaluation and ensures the benchmark’s overall quality.
As we analyzed the 118 samples excluded during the ELITE benchmark filtering process (Not From ELITE), we observed that 52.54% of them resulted in responses that were merely descriptive of the input image. This highlights the need for a toxicity score to distinguish genuinely harmful content, ensuring a more accurate and meaningful safety evaluation.
Table 8: Human safety labels for From ELITE and Not From ELITE samples.
Majority Vote From ELITE Not From ELITE
Unsafe 67.27% 11.86%
Safe 32.73% 88.14%
@@ 7 Limitations
The ELITE evaluator aims to provide an accurate and reliable safety evaluation for vision-language models. However, as a rubric-based approach, its evaluation performance may vary depending on the capabilities of VLMs, which we acknowledge as a potential limitation.
Moreover, although the ELITE benchmark is carefully constructed to cover a broad taxonomy and include a wide range of diverse image-text pairs, it may still miss certain cases that can elicit harmful responses. For instance, more complex scenarios such as multi-turn jailbreak techniques are not yet incorporated, suggesting room for future expansion.
@@ 8 Conclusion
In this work, we introduce the ELITE evaluator, a rubric-based method enhanced with toxicity scoring to provide more precise and human-aligned safety assessments. This evaluator effectively overcomes the limitations of existing evaluation methods, which frequently struggle to separate truly harmful outputs from harmless yet uninformative responses, such as those that merely describe the input image. Alongside the evaluator, we construct the ELITE benchmark, a high-quality dataset of 4,587 image-text pairs curated through the ELITE evaluator. By filtering out ambiguous image-text pairs and incorporating in-house generated image-text pairs that cover all image-text pairs—including safe-safe combinations—we ensure the benchmark supports broad and robust safety evaluation. Our experimental results demonstrate that the ELITE evaluator aligns more closely with human judgment than existing methods and that the ELITE benchmark exposes harmful responses more effectively than prior benchmarks. We hope this work facilitates future research in multimodal safety and contributes toward the development of more reliable, trustworthy VLMs.
Acknowledgments
This research was partly supported by the MSIT(Ministry of Science and ICT), Korea, under the ITRC(Information Technology Research Center) support program(IITP-2024-RS-2023-00258649, 50%) supervised by the IITP(Institute for Information & Communications Technology Planning & Evaluation), partly supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2025-00562437, 30%), and partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2022-00155911, Artificial Intelligence Convergence Innovation Human Resources Development (Kyung Hee University), 10%), and by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2022-00143524, Development of Fundamental Technology and Integrated Solution for Next-Generation Automatic Artificial Intelligence System, 10%)
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. In this work, we introduce a benchmark to evaluate the safety of VLMs. Given its nature, the benchmark contains potentially offensive samples, which may raise safety concerns. We affirm that all data used in this study will not be utilized for purposes other than research. Our research aims to focus on the safety challenges of VLMs and to facilitate future research on their safety alignment to prevent harmful responses
@@ A.1 Benchmark using the ELITE Evaluator Score
Table 9: ELITE evaluator score of various VLMs across taxonomies. The upper group in the table represents proprietary models, and the lower group represents open-source models. Highlight the most vulnerable model in bold and the second-most vulnerable with an underline. * denotes the model used for filtering.
@@ B.2 Previous Benchmarks
SIUO (Wang et al., 2025) 167 Human Human, GPT-4V
@@ B.5 Model Cards
Qwen2-VL-7B (Wang et al., 2024) 7B Custom ViT with 600M params Qwen2-7B